# Reconstructing Galaxy Cluster Mass Maps using Score-based Generative Modeling

### v1

@@ 3.2 Conditional Score-based Models
Our score model transforms a sample from a Gaussian random field into a realization of the gas and dark matter map of a galaxy cluster, which effectively samples from the gas and dark matter posterior distribution. Figure 3 describes the training process of a score-based model using SMLD (e.g. Song et al., 2021) using stochastic differential equations (SDEs). We provide the forward SDE to gradually transform the gas and dark matter maps into random noise, and the diffusion model learns the inverse SDE that maps the noise back to the gas and dark matter maps. In particular, the model learns to sample from the joint posterior of gas and dark matter maps of clusters conditioned on the SZ and X-ray inputs. This section discusses the statistical framework behind the sampling techniques for the posterior, and the training process for a score-based diffusion model.
@@
$d\bm{x}=\bm{f}(\bm{x},t)dt+g(t)d\bm{w},$ (8)
also shown at the top of Figure 3. Here, $\bm{x}=(\text{Gas},\text{DM})$ are the gas and dark matter maps, the $\bm{f}$ and $g$ components are the drift and diffusion coefficients, and $\bm{w}$ is the standard Wiener process. The forward SDE maps $\bm{x}(t=0)$ sampled from the data distribution to $\bm{x}(t=T)$ sampled from a Gaussian random field.
The model learns to sample the data distribution $p_{t}(x)$ by solving the reverse-time SDE (e.g. Anderson, 1982),
$d\bm{x}=[\bm{f}(\bm{x},t)-g^{2}(t)\nabla_{\bm{x}}\log p_{t}(\bm{x})]dt+g(t)d\bm{\bar{w}},$ (9)
which inverts the forward-time transformation by mapping $\bm{x}(t=T)$ to $\bm{x}(t=0)$, also shown at the bottom of Figure 3. During training, the model estimates the score ($\nabla_{\bm{x}}\log p_{t}(\bm{x})$), or the gradient with respect to the data of the log probability density at some time $t$, by parameterizing it with a deep network $s_{\theta}(\bm{x},t)$.
The actual training consists of solving for the optimal $\theta^{*}$ via denoising score matching (e.g. Vincent, 2011; Song et al., 2021) by optimizing for the following objective:
@@
Here, we are finding the best $\theta^{*}$ such that our deep network $s_{\theta}$ best reconstructs, in the L2-norm, the true score of the distribution in expectation across all $t\in[0,T]$ and all $\bm{x}(t)$. In practice, $\nabla_{\bm{x}(t)}\log p_{t}(\bm{x}(t))$ is evaluated as the numerical derivative between subsequent noising steps.
[…]
$\begin{split}d\bm{x}&=[\bm{f}(\bm{x},t)-g^{2}(t)(\nabla_{\bm{x}}\log p_{t}(\bm{x})+\nabla_{\bm{x}}\log p_{t}(\bm{y}|\bm{x}))]dt\\ &+g(t)d\bm{\bar{w}},\end{split}$ (11)
where we learn an additional gradient of the conditional forward process, $\nabla_{\bm{x}}\log p_{t}(\bm{y}|\bm{x})$, as part of our parameterized network. This will be the actual reverse SDE we use in training.
Our diffusion model uses the Noise Conditional Score Network (NCSN++) architecture and a Variance-Exploding SDE sampling process (VESDE, as described in appendix B of Song et al., 2021) to learn the conditional score of the distribution. The NSCN++ is a variant of RefineNet (Lin et al., 2016), a U-Net variant with instance normalization and dilated convolutions (c.f.  Song & Ermon, 2019, appendix A). For training, the model has 2 residual blocks per resolution, and $\sigma_{max}=1$, the largest intensity of our images, and $\sigma_{min}=0.001$, which sets the scale of the noise in our images. We use the implementation given by Alexendre Adam’s repository (e.g. Adam et al., 2022; Legin et al., 2023) that can be found at this link: https://github.com/AlexandreAdam/score_models.
Figure 4: Selected model predictions of gas and dark matter maps (rightmost 2 columns) conditioned on SZ and X-ray inputs (leftmost 2 columns) along with ground truth gas and dark matter maps (middle 2 columns). The predicted maps are averages of 50 samples from the learned joint posterior. Across large and small mass clusters (different rows), we see accurate reconstruction for large and small-scale structures.
@@ 4 Results
To validate generated results from our score model, we apply a series of performance metrics to our model with an independent testing set of mock observations. We present sample generated gas and dark matter maps along with their corresponding sampling convergence and fractional standard error (Section 4.1). We then compute the radial density profiles (Section 4.2) and bias and cross correlation coefficients (Section 4.3) and demonstrate that our model is able to generate maps that reconstruct these quantities with high accuracy.
Figure 5: Sample variance of mean pixel gas predictions (purple) and dark matter (orange). Since each starting sample is a different noise realization, the output gas and dark matter maps will be slightly different as well: the sample variance is hence a way of assessing whether the mean of these output images converged. As we increase the number of samples, the variance decreases as we converge to the average value of the posterior distribution: at around 40 samples, the variance goes below the 95th percentile of a 2-sample variance (dotted lines). We thus select 50 samples as a safe threshold when computing our average predictions.
@@ 4.1 Mass Maps
The diffusion model is trained on clusters of sizes $1.00\times 10^{14}\leq M_{200a}/M_{\odot}\leq 4.82\times 10^{15}$, but we will only test our model on clusters within the range $2.00\times 10^{14}\leq M_{200a}/M_{\odot}\leq 1.00\times 10^{15}$ since we lack sufficient data to constrain cluster predictions at these extreme high and low mass tails. Previous deep-learning based estimates of cluster masses (e.g. Ho et al., 2019; Ntampaka et al., 2019; Yan et al., 2020) broadly observed a ‘mean-reversion’ effect, wherein the masses of small clusters were overpredicted and those of large clusters were underpredicted. Such effects can be resolved with more training data and broader mass priors. As a result, we exclude these extremities and reserve their analysis for future work. For our test mass region, we will denote clusters within $2.00\times 10^{14}\leq M_{200a}/M_{\odot}<6.00\times 10^{14}$ as low mass clusters and $6.00\times 10^{14}\leq M_{200a}/M_{\odot}\leq 1.00\times 10^{15}$ as high mass clusters, roughly divided evenly in log mass space.
The diffusion model samples from the learned posterior by first sampling from a Gaussian random field, and mapping that to a realization of gas and dark matter densities. Since the sampling process is probabilistic, the sample variance is used to determine the number of samples we need from the posterior to produce a converged mean prediction. In Figure 5, we compute the 95 percentile convergence of the sample variance averaged over all pixels, which is around 40 samples for both the gas and dark matter. We will use 50 samples to compute our results.
Figure 4 shows sample galaxy cluster predictions from our diffusion model, with each row being a sample from a different mass bin. The samples are selected to provide a representative overview of the performance of our model on clusters of different shapes and sizes ranging across our full test mass range. The first 4 columns represent the ground truth simulation maps of SZ, X-ray, gas, and dark matter, respectively. Multiple realizations of the gas and dark matter maps are generated to compute statistical quantities: the final 2 columns represent the generated 50-sample average of the gas and dark matter maps conditioned on the SZ and X-ray inputs. Visually, our model is able to pick up on both the large-scale and small-scale features of the clusters across different mass bins (rows). We further justify the accuracy of our model through statistical metrics.
Figure 6: PDF of standardized pixel-wise fractional mass error (Equation 12) of generated gas (purple) and dark matter (orange) maps. The distribution is centered at zero with a smaller-than-normal spread, indicating that our model is under-confident but, on average, unbiased on the pixel level.
To understand the bias on a pixel level, we validate our gas and dark matter predictions using a histogram of pixel-level standardized errors $\epsilon$ given by the difference between the predicted $S_{\mathrm{gen}}$ and the simulated $S_{\mathrm{sim}}$ divided by the variance of the predictions $\sigma_{\mathrm{gen}}$,
$\epsilon_{\mathrm{gen}}[i,j]=\frac{S_{\mathrm{gen}}[i,j]-S_{\mathrm{sim}}[i,j]}{\sigma_{\mathrm{gen}}[i,j]}.$ (12)
@@
Figure 7: Radial density profiles for dark matter (left) and gas (right). Large cluster profiles are shown on the top, and small cluster profiles are shown on the bottom, and in each panel we graph the generated maps (magenta) and simulated maps (cyan) with the 16th to 84th percentiles shaded. Under each curve, we show the fractional difference between the generated and simulated maps in red, which shows agreement to within 5% across all radial distances. We see that the model slightly overpredicts (positive fractional difference) the cluster density until  3 Mpc, where the fractional difference begins to decline and eventually underpredicts (negative fractional difference). This trend follows the expected behavior to match the zero-density boundary condition of our mass maps.
Figure 6 shows the probability density (normalized histogram) of the error for the gas and dark matter maps, which is approximately normal about zero with a smaller-than-unity spread. As expected, for both mass maps the errors are centered around zero, indicating zero bias during the prediction pipeline of our model on average. However, the spread is much smaller than a standard normal, as 99.5% of the data is contained within ±$\pm 1\sigma$. Since $\epsilon_{\mathrm{gen}}$ is computed by dividing by $\sigma_{\mathrm{gen}}$, the pixel-wise variance of the dark matter and gas maps is actually larger than our posterior variance, indicating that the model is under-confident in its predictions. This metric shows that we did not overfit our model to the training dataset and that the diffusion sampling process does not exhibit mode collapse.
Figure 8: Bias coefficients (blue) between the simulation and generated outputs (Equation 13, for both high mass clusters (top) and low mass clusters (bottom). The model achieves close-to-unity bias (black) across all scales.
@@ 4.2 Density Profiles
To validate the performance of mass reconstruction radially, we compute the average density profile for both the generated and simulated gas and dark matter maps, binning the pixels radially. Figure 7 shows these profiles on the top panel, along with the associated fractional reconstruction error in the bottom panel. Overall we match both the mean and the 68% spread in our reconstructed mass maps with the ground truth mass distribution, with a 1-2% average fractional error across all clusters in the central $r<1$ Mpc region, with a slightly larger fractional error but still within 5% at the edge of the cluster in regions $r\geq 1$ Mpc. Matching both the mean trend and spread is equally important, as there is natural scatter from the simulation profiles from the various cluster sizes: the fact that our model is able to reproduce this uncertainty is a strong indication that it is able to distinguish between clusters of different sizes.
We further demonstrate that our model tries to preserve the overall mass: the average total mass fractional error for the gas maps is $0.005^{+0.028}_{-0.028}$, while that of the dark matter is $0.008^{+0.023}_{-0.024}$. Thus, the slight over-prediction near the center of the cluster is balanced by the slight under-prediction past the virial radius, where the model is also trying to match the zero boundary condition near the edge of the images. The increase in the magnitude of the error beyond the virial radius is an artifact of the denominator in the fractional error: because there is minimal mass near the edge, the residual (numerator) becomes much larger than the denominator.
Our results improve on the error of those of Figure 6 in de Andres et al. (2024), where they compute the average total mass fractional error of generated density maps from SZ, X-ray and star observations using a U-Net model to $-0.00^{+0.23}_{-0.18}$. Our model leverages the sampling power of diffusion models to converge to a much more accurate gas and dark matter map rather than a single point estimate as used in other image-to-image models.

### v2

@@ 3.2 Conditional Score-based Models
Our score model transforms a sample from a Gaussian random field into a realization of the gas and dark matter map of a galaxy cluster, which effectively samples from the gas and dark matter posterior distribution. Figure 3 describes the sampling process of a score-based model using stochastic differential equations (SDEs) (e.g. Song et al., 2021). We provide the forward SDE to gradually transform the gas and dark matter maps into random noise, and the diffusion model learns the reverse SDE that maps the noise back to the gas and dark matter maps. In particular, the model learns to sample from the joint posterior of gas and dark matter maps of clusters conditioned on the SZ and X-ray inputs. This section discusses the statistical framework behind the sampling techniques for the posterior, and the training process for a score-based diffusion model.
@@
$d\bm{x}=\bm{f}(\bm{x}(t),t)dt+g(t)d\bm{w},$ (8)
also shown at the top of Figure 3. Here, $\bm{x}=(\text{Gas},\text{DM})$ are the gas and dark matter maps, which are a function of time $t$ that defines the noise level of the process. $\bm{f}$ is the drift component, which describes the deterministic force that pushes on $\bm{x}$: for our case, this will be 0, as $\bm{x}$ eventually evolves into purely random noise. $g$ is the diffusion component, which describes the scale of noise added to the system. $\bm{w}$ is the standard Wiener process, or Brownian motion. The forward SDE maps $\bm{x}(t=0)$ sampled from the data distribution to $\bm{x}(t=T)$ sampled from a Gaussian random field by stochastically integrating Equation 8 using standard SDE methods such as Euler-Maruyama (e.g. Maruyama, 1955; Song et al., 2021).
The model learns to sample the data distribution $p_{t}(\bm{x})$ by solving the reverse-time SDE (e.g. Anderson, 1982)
$d\bm{x}=[\bm{f}(\bm{x}(t),t)-g^{2}(t)\nabla_{\bm{x}}\log p_{t}(\bm{x})]dt+g(t)d\bm{\bar{w}},$ (9)
which inverts the forward-time transformation by mapping $\bm{x}(t=T)$ to $\bm{x}(t=0)$, also shown at the bottom of Figure 3. This equation requires knowing the score of the data distribution $\nabla_{\bm{x}}\log p_{t}(\bm{x})$, or the gradient with respect to the data of the log probability density at some noise level $\sigma(t)$. We estimate the score using machine learning by parameterizing it with a deep network $s_{\theta}(\bm{x}(t),t)$.
Learning the distribution of $\nabla_{\bm{x}}\log p_{t}(\bm{x})$ directly is complex, since we do not have access to the true data distribution $p(\bm{x})$, and thus we can only reply on supervised signals derived from known conditionals: Song et al. (2021) uses a technique that introduces noise to the data distribution, and solves for the conditional noisy distribution instead. For some parameterization of the noise $\sigma(t)$, we can reparameterize $\bm{x}(t)$ as a noisy data point $\bm{\tilde{x}}=\bm{x}+\sigma\xi$, where $\xi\sim N(0,1)$ is a normal distribution or any known noise distribution. Then, $p(\bm{\tilde{x}}|\bm{x})$ is analytically a Gaussian, and thus the score is also analytically known:
$\nabla_{\bm{\tilde{x}}}\log p_{\sigma}(\bm{\tilde{x}}|\bm{x})=-\frac{\bm{\tilde{x}}-\bm{x}}{\sigma^{2}}.$ (10)
Using a key result from the denoising score matching procedure (e.g. Vincent, 2011; Song et al., 2021), optimizing the score matching relative to $\nabla_{\bm{\tilde{x}}}\log p_{\sigma}(\bm{\tilde{x}}|\bm{x})$ implicitly optimizes the score matching to the true marginal score $\nabla_{\bm{\tilde{x}}}\log p_{\sigma}(\bm{\tilde{x}})$ up to a constant. In other words, instead of solving a very hard optimization problem that uses the objective
$\underset{p_{\sigma}(\bm{\tilde{x}})}{\mathbb{E}}\norm{s_{\theta}(\bm{\tilde{x}},\sigma)-\nabla_{\bm{\tilde{x}}}\log p_{\sigma}(\bm{\tilde{x}})}_{2}^{2},$ (11)
we utilize a known form of the conditional score in equation 10 and optimize the alternative objective
$\underset{p(\bm{x})p(\bm{\tilde{x}}|\bm{x})}{\mathbb{E}}\norm{s_{\theta}(\bm{\tilde{x}},\sigma)+\frac{\bm{\tilde{x}}-\bm{x}}{\sigma^{2}}}_{2}^{2},$ (12)
which effectively minimizes the normalized difference between the noised and unnoised maps. This method cleverly utilizes the conditional distribution only during training as a surrogate to optimize the marginal score that we want to estimate. In practice, the expectation is approximated by the samples of our data, and the optimization process produces the best $\theta^{*}$ that best matches the score. The trained model $s_{\theta^{*}}(\bm{\tilde{x}},\sigma)$ is then used as an estimate of the score term $\nabla_{\bm{x}}\log p_{t}(\bm{x})$ in equation 9. Note that $\sigma$ and $t$ are used interchangeably here, and in particular we describe the noising process in SDEs with the independent variable $t$, but during score matching we describe the noise as $\sigma$.
@@
$\begin{split}d\bm{x}&=[\bm{f}(\bm{x},t)-g^{2}(t)\nabla_{\bm{x}}\log p_{t}(\bm{x}|\bm{y})]dt+g(t)d\bm{\bar{w}}.\end{split}$ (13)
In the case where we analytically understand the likelihood of our samples $p(\bm{y}|\bm{x})$, we would use Bayes rule to separate the prior from the likelihood in the score, namely $\nabla_{\bm{x}}\log p_{t}(\bm{x}|\bm{y})=\nabla_{\bm{x}}\log p_{t}(\bm{x})+\nabla_{\bm{x}}\log p(\bm{y}|\bm{x})$, and only use machine learning to estimate the first term and analytically evaluate the second. However, since it is difficult to express the relationship between the SZ and X-ray maps to the Gas and DM maps analytically, we will parameterize the full score $s_{\theta}(\bm{\tilde{x}},t,\bm{y})$ using a model that takes in $\bm{y}$ as additional channels. The corresponding objective is then modified to perform an expectation over all $(\bm{x},\bm{y})$ data pairs:
$\underset{p(\bm{x,y})p(\bm{\tilde{x}}|\bm{x})}{\mathbb{E}}\norm{s_{\theta}(\bm{\tilde{x}},\sigma,\bm{y})+\frac{\bm{\tilde{x}}-\bm{x}}{\sigma^{2}}}_{2}^{2},$ (14)
and once trained this conditional network is likewise substituted into the score term in equation 13.
Finally, the noising procedure that we will use is described by the exponential schedule
$\sigma(t)=\sigma_{min}*\left[\frac{\sigma_{max}}{\sigma_{min}}\right]^{t/T},$ (15)
where the variance at time $t=0$ is $\sigma_{min}$, defined to be the scale of noise in the data distribution, and the variance at time $t=T$ is $\sigma_{max}$, which sets the scale of the noise in the Gaussian random field. In general, we want to set $\sigma_{min}$ to be a reasonable lower bound for the noise we find and want to reconstruct in the data, and $\sigma_{max}$ to be comparable to the normalized pixel values of the images. A balance is to be found for the range of these variance bounds: a noising procedure with a small change in $\sigma$ may not learn the small-scale structures and desired accuracy in the data distribution, while a large change in $\sigma$ may not be able to learn the data distribution at all.
Our diffusion model uses the Noise Conditional Score Network (NCSN++) architecture and a Variance-Exploding SDE sampling process (VESDE, as described in appendix B of Song et al., 2021) to learn the conditional score of the distribution. The NSCN++ is a variant of RefineNet (Lin et al., 2016), a U-Net variant with instance normalization and dilated convolutions (c.f.  Song & Ermon, 2019, appendix A). For training, the model has 2 residual blocks per resolution, $\sigma_{max}=1$, and $\sigma_{min}=0.001$. For sampling, the variance-exploding procedure has a drift term of $\bm{f}=0$ and a diffusion term $g=\sqrt{d\sigma^{2}/dt}$. We use the implementation given by Alexendre Adam’s repository (e.g. Adam et al., 2022; Legin et al., 2023) that can be found at this link: https://github.com/AlexandreAdam/score_models.
@@ 4 Results
To validate generated results from our score model, we apply a series of performance metrics to our model with an independent testing set of mock observations. We present sample generated gas and dark matter maps along with their corresponding sampling convergence (Section 4.1). We then compute the radial density profiles and pixel-wise fractional error histograms (Section 4.2) and bias and cross correlation coefficients (Section 4.3), and demonstrate that our model is able to generate maps that reconstruct these quantities with high accuracy.
Figure 4: Sample variance of mean pixel gas predictions (purple) and dark matter (orange) as a function of the number of samples. Since each starting sample is a different noise realization, the output gas and dark matter maps will be slightly different as well: the sample variance is hence a way of assessing whether the mean of these output images converged. As we increase the number of samples, the variance decreases as we converge to the mean value of the posterior distribution: at around five samples, the variance undergoes one e-folding of a two-sample variance (dotted lines). We thus select ten samples as a safe threshold when computing our average predictions.
@@ 4.1 Mass Maps
Figure 5: Selected model predictions of gas and dark matter maps (rightmost 2 columns) conditioned on SZ and X-ray inputs (leftmost 2 columns) along with ground truth gas and dark matter maps (middle 2 columns). For each test cluster, we show the linear maps on the top row and the log-transformed maps on the bottom row. The predicted maps are averages of 10 samples from the learned joint posterior. While our models are trained on data in linear scale, the images are also displayed in logarithmic scale to enhance dynamic range and highlight substructure. Across large and small mass clusters, we see accurate reconstruction for large and small-scale structures, particularly noticeable when we log-transform the generated images.
@@
Figure 6: Radial density profiles for dark matter (left) and gas (right). Profiles of high mass clusters ($6\times 10^{14}\leq M_{200m}/M_{\odot}\leq 1\times 10^{15}$) are shown on the top, and those of low mass clusters ($2\times 10^{14}\leq M_{200m}/M_{\odot}<6\times 10^{14}$) are shown on the bottom. In each panel we graph the generated maps (magenta) and simulated maps (cyan) with the 16th to 84th percentiles shaded. The ranges for $R_{200m}$ (shaded gray) in Mpc are $2.60\leq R_{200m}\leq 3.06$ for large clusters and $1.81\leq R_{200m}\leq 2.60$ for small clusters. Under each curve, we show the fractional difference between the generated and simulated maps in red. For the gas, we see agreement to within 5% for all scales for both high and low mass clusters, with a slightly larger fractional error for $r>1$ Mpc. For the dark matter, the fractional error spread increases to around 10% at the smallest spatial scales.
The diffusion model is trained on clusters of sizes $1.00\times 10^{14}\leq M_{200m}/M_{\odot}\leq 4.82\times 10^{15}$, but we will only test our model on clusters within the range $2.00\times 10^{14}\leq M_{200m}/M_{\odot}\leq 1.00\times 10^{15}$. Previous deep-learning based estimates of cluster masses (e.g. Ho et al., 2019; Ntampaka et al., 2019; Yan et al., 2020) broadly observed a ‘mean-reversion’ effect, wherein the masses of small clusters were overpredicted and those of large clusters were underpredicted. Such effects can be resolved with more training data and broader mass priors. As a result, we exclude these extremities and reserve their analysis for future work. For our test mass region, we will denote clusters within $2.00\times 10^{14}\leq M_{200m}/M_{\odot}<6.00\times 10^{14}$ as low mass clusters and $6.00\times 10^{14}\leq M_{200m}/M_{\odot}\leq 1.00\times 10^{15}$ as high mass clusters, roughly divided evenly in log mass space. The ranges for $R_{200m}$ in Mpc are $2.60\leq R_{200m}\leq 3.06$ for large clusters and $1.81\leq R_{200m}\leq 2.60$ for small clusters.
The diffusion model samples from the learned posterior by first sampling from a Gaussian random field, and mapping that to a realization of gas and dark matter densities using the reverse SDE and the specified noising schedule. Since the sampling process is probabilistic, the sample variance is used to determine the number of samples we need from the posterior to produce a converged mean prediction. In Figure 4, we compute the sample variance averaged over all pixels as a function of the number of samples, and we observe that one e-folding of a two-sample variance occurs at around five samples for both the gas and dark matter. We will use ten samples, each of which is generated from a random Gaussian random field, to compute mean predictions in the following results. The samples are generated via the noising equation 15 with $1000$ discrete time steps.
Figure 5 shows three sample galaxy cluster predictions from our diffusion model, ranging across our full test mass range. Each pair of rows is a test cluster, with linear maps on the top row and the log-transformed maps on the bottom row. For the log-transformed maps, the dark matter and gas densities are clipped to a mean density of 1 before we apply the log-transform, and the SZ and X-ray inputs are clipped to a corresponding reasonable lower bound such that the log-transform shows the major substructures. While our models are trained on data in linear scale, the images are also displayed in logarithmic scale to enhance dynamic range and highlight substructure. The first 4 columns represent the ground truth simulation maps of SZ, X-ray, gas, and dark matter, respectively. Multiple realizations of the gas and dark matter maps are generated to compute statistical quantities: the final 2 columns represent the generated 10-sample average of the gas and dark matter maps conditioned on the SZ and X-ray inputs. Visually, our model is able to pick up on both the large-scale and small-scale features of the clusters across different mass bins (rows), which are especially noticeable in the log-transformed images. While the substructure signals in the SZ and X-ray may not be large in the simulation images, they are still present, and we find that the machine learning model is able to pick on these small signals in those regions and learn a complex internal model of the galaxy cluster to produce the substructures in the output gas and dark matter images.
@@ 4.2 Density Profiles
Figure 7: Fractional error distribution for dark matter (orange) and gas (purple), for pixels inside $R_{200m}$ (solid line) and outside $R_{200m}$ (dashed line). We see a smaller spread in error for pixels inside $R_{200m}$, which indicates that our model reconstructs pixels within $R_{200m}$ better than those outside. This matches the error curve in the radial profiles shown in Figure 6, where we see low error near the center that increases in both bias and variance as we increase the radius.
To validate the performance of mass reconstruction radially, we compute the average density profile for both the generated and simulated gas and dark matter maps, binning the pixels radially. Figure 6 shows these profiles on the top panel, along with the associated fractional reconstruction error in the bottom panel. The simulation average radial profiles are shown in magneta, and the generated average radial profiles are shown in cyan, with the ranges of $R_{200m}$ shaded vertically in gray. Below each plot we show the fractional error in red. Overall we match both the mean and the 68% spread in our reconstructed mass maps with the ground truth mass distribution, with a 1-2% average fractional error across all clusters in the central $r<1$ Mpc region, with a slightly larger fractional error at the edge of the cluster in regions $r\geq 1$ Mpc. We also find that the standard deviation of the fractional error is the same order of magnitude as the bias. Matching both the mean trend and spread is equally important, as there is natural scatter from the simulation profiles from the various cluster sizes: the fact that our model is able to reproduce this uncertainty is a strong indication that it is able to distinguish between clusters of different sizes.
We further demonstrate that our model tries to preserve the overall mass: the average total mass fractional error for the gas maps is $0.0023^{+0.0010}_{-0.0018}$, while that of the dark matter is $-0.0079^{+0.0070}_{-0.0044}$. Overall, we find that for both gas and dark matter, the reconstructed large clusters have a slight under-bias at $r>1$ Mpc where the model tries to match the zero boundary condition near the edges of the maps. For small clusters, the model slightly overpredicts near the edges, where increase in the magnitude of the error beyond the virial radius is an artifact of the denominator in the fractional error: because there is minimal mass near the edge, the residual (numerator) becomes much larger than the denominator. Finally, in Figure 7 we show the pixel-wise fractional error distribution of the reconstructed images, split into pixels within and outside of $R_{200m}$. We see that the model more accurately reconstructs pixels within $R_{200m}$, while those outside $R_{200m}$ have a much flatter fractional error distribution.
Figure 8: Bias coefficients (blue) between the simulation and generated outputs (Equation 16). Coefficients for high mass clusters ($6\times 10^{14}\leq M_{200m}/M_{\odot}\leq 1\times 10^{15}$) are shown on the top, and those of low mass clusters ($2\times 10^{14}\leq M_{200m}/M_{\odot}<6\times 10^{14}$) are shown on the bottom. The model achieves close-to-unity bias (black) across all scales.
Our results improve on the error of those of Table 4 in de Andres et al. (2024), where they compute the average total mass fractional error of generated density maps from SZ, X-ray and star observations using a U-Net model to be $-0.015^{+0.028}_{-0.026}$ for clusters of size $14.6314.02\leq\log M_{200}/h^{-1}M_{\odot}<$ 14.63, and $-0.024^{+0.019}_{-0.017}$ for clusters of size $15.3814.63\leq\log M_{200}/h^{-1}M_{\odot}<$ 15.38. Our model leverages the sampling power of diffusion models to converge to a much more accurate gas and dark matter map rather than a single point estimate as used in other image-to-image models.